OpenStudy (anonymous):

if the integral from 3 to 6 of7f(x)-2 =7 what does f(x) equal on the same interval.....please help me

OpenStudy (loser66):

where are you stuck?

OpenStudy (anonymous):

I guess I'm not sure what what the integral of 3 to 6 of 7*(f(x)-2dx=7 really even means without the actual function

OpenStudy (loser66):

\[\int_3^6 (f(x)-2)dx=7\]

OpenStudy (loser66):

but I am not sure whether it is f(x)-2 or f(x-2)

OpenStudy (anonymous):

you had it right in your first reply f(x) -2

OpenStudy (loser66):

ok, so why are you stuck? it's quite easy

OpenStudy (anonymous):

that's the thing it seems like it should be really easy but i really don't have a method for computing what f(x) is i guess i've never seen an integration question posed like this

OpenStudy (loser66):

\[\int_3^6 (f(x)-2)dx = \int_3^6 f(x)dx -\int_3^6 2dx =7\] \[\int_3^6 (f(x)=7+\int_3^6 2dx = 7 + (2*6-2*3)= 13\]

OpenStudy (anonymous):

its 7*f(x) -2 but your answer actually gives me what I need thanx

OpenStudy (loser66):

so, just divide by 7

OpenStudy (anonymous):

right thank you I appreciate it I think forgot about those old properties of integrals